OpenStudy (anonymous):

HELP PLEASE!!!!!!!!!! O.o Solve the following system of equations graphically. y = 2x - 3 x + y = 3 What is the solution set? A. {(2, 1)} B. {(1, 2)} C. {(-1, -2)} D. {(-2, -1)}

OpenStudy (solomonzelman):

it is set as, for a substitution: Substitute what the y is equal to in a first equation, INTO a second equation (for y). (read this sentence carefully)

jimthompson5910 (jim_thompson5910):

You can use a graphing calculator like this https://www.desmos.com/calculator/ if you don't have a graphing calculator

jimthompson5910 (jim_thompson5910):

graph each equation in each line and the point of intersection is your solution

OpenStudy (solomonzelman):

oh, graphically! I read algebraically for some reason-:(

OpenStudy (solomonzelman):

btw, I also use desmos:) ....

OpenStudy (triciaal):

graph each line on the same grid and look where they intersect

OpenStudy (triciaal):

@Krystin_56 do you know how to graph without electronics?

OpenStudy (anonymous):

sorta @triciaal

OpenStudy (triciaal):

write equation in the form y = mx + b to get a line you need 2 points the easiest to use when x = 0 what is y and when y = 0 what is x join the 2 points do the same for the other line solution is the intersection of the 2 lines

OpenStudy (triciaal):

y = 2x - 3 when x = 0 y = -3 when y = 0 2x -3 = 0 so x = 3/2 ((0, -3) and (3/2, 0) x + y = 3 when x = 0 y = 3 when y = 0 x = 3 (0, 3) and (3,0)

OpenStudy (triciaal):